Divide rational expression, and if possible, simplify.
a^2 - b^2 /a^2 - 4ab + 4b^2 ÷ a^2 - 3ab + 2b^2/ a - 2b
I know to rewrite as multiplication of reciprocal next step
= a^2 - b^2/a^2 - 4ab + 4b^2 ÷ a - 2b/ a^2 - 3ab + 2b^2

Factor next step and cancel common factors
a^2 - b^2/ a^2 - 4ab+4b^2 ÷ a - 2b /a^2 -3ab + 2b^2
Not sure what’s that next step for a solution.
Thank you for any assistance.






Answer by london maths tutor(243)

Factorise the following:
a^2 - b^2 = (a+b)(a-b)
a^2 - 3ab + 2b^2 = (a-2b)(a-b)
Put these factorised brackets into the equation and change the question to multiplication.
you should get:
[(a+b)(a-b)/(a^2-4ab+4b^2)] * [(a-2b)/(a-2b)(a-b)]
you can cancel out (a-2b) and (a-b)
The remaining fraction becomes:
[(a+b)/(a^2-4ab+4b^2)]